Light of wavelength 730nm falls on two slits and produces an interference pattern in which the third-order bright red fringe is 36mm from the central fringe on a screen 1.8m away.

What is the separation of the two slits?
Answer #1

we have:

y =\frac{m\lambda L}{d}

36\times 10^{-3} =\frac{3(730\times 10^{-9}) 1.8}{d}

solving for d we have:

d = 1.095\times 10^{-4}m